# Closing the Data-Efficiency Gap Between Autoregressive and Masked Diffusion LLMs

### v2

@@ 3 Datasets and experimental setups
We first show that knowledge injection by fine-tuning in arLLMs heavily relies on paraphrases. This is known in previous studies (berglund2023reversal; zhu24b; lin2024delving; guo-etal-2024-mitigating; golovneva2024reverse). We demonstrate this observation on three datasets to set baselines for comparison with dLLM and our novel paradigm in the following sections.
@@
We fine-tune arLLMs on dataset samples using the pre-training format (i.e., without a chat template). Without paraphrases, backward accuracy on the NameDescription and Biography datasets is close to 0, while the forward accuracy of NameDescription N2D and Biography does not completely fail but is still poor (Table 1). Adding same-order paraphrases drastically raises forward accuracy close to 1, while backward accuracy remains close to 0. Paraphrases do not help backward accuracy in NameDescription and Biography datasets because the construction of these datasets does not change the semantic order of the sentences. The trend is similar in the Wiki dataset (Table 1 and 2). While the same-order paraphrases significantly increase forward accuracy, they only mildly increase backward accuracy. Using permute-order paraphrases increases both forward and backward accuracy, and the gap between them is smaller. Note that, due to the naturalness of this dataset, pre–fine-tuning accuracies are not as close to zero as in the other datasets (Table 1 and 2); nonetheless, they are sufficiently low to demonstrate the effectiveness of fine-tuning.
These results suggest that, in arLLM fine-tuning, paraphrases significantly improve QA accuracy, but help backward questions only when the paraphrases change the information order in the original text to be more aligned with the backward style. Note that the accuracy difference between fine-tuning with paraphrases and without paraphrases is not due to different training steps; in both cases, we train the models with sufficiently large epoch numbers; the reported accuracy is from the best checkpoints during the training (Figure 2, Appendix 6).
5 dLLM knowledge injection
@@
We investigate the data efficiency of dLLMs for knowledge injection by fine-tuning, focusing on whether they require paraphrases to successfully handle both forward and backward QAs. We follow the original pretraining protocol (llada) to fine-tune LLaDA-8B-Instruct on the dataset samples using the loss defined in Eq. 1. On three datasets, the accuracy difference between fine-tuning with and without paraphrases is much smaller in the dLLM than in the arLLM (Table 1): dLLM without paraphrases can already achieve decent and comparable accuracies on both forward and backward questions; fine-tuning with paraphrases can further increase the accuracy by a small amount. These results together suggest that dLLM has superior data efficiency and is free of the reversal curse in the post-training phase. By plotting the testing accuracy across the training steps (Figure 2), we observe that arLLM fine-tuned without paraphrases improves QA accuracy only in the beginning of training, then quickly decreases, indicating overfitting. The dLLM without paraphrases, on the other hand, does not show signs of overfitting. This finding echoes what has been found in comparing arLLMs and dLLMs in the pre-training phase (prabhudesai2025diffusion; ni2025difflm).
One may expect that fine-tuning dLLM converges slower than arLLM, because learning any-order factorization requires seeing more than one way of the factorizations (i.e., samples masked in different ways) (xue2025any; kim2025train). However, we found that dLLM converges at least as fast as arLLM (Figure 2, Table 4, Appendix Table 5); in the Biography dataset, dLLM even converges faster than arLLM. This indicates that dLLM does not trade better data efficiency and performance for more training compute; it requires the same or less training compute and fewer training samples but achieves better downstream performance.
@@
Figure 2: Training dynamics of arLLM (Llama 8B), dLLM (Llada), and masked arLLM (Llama 8B). For the NameDescription dataset, forward and backward accuracy are the average of N2D and D2N types. Paraphrases used in the Wiki dataset are the same-order paraphrases set. Due to the randomness of sampling the masks, we average across 4 random seed for the dLLM and masked arLLM on NameDescription and Biography Datasets. Curves for each seed are shown in Appendix Figure 10-11.
6 Masked Fine-tuning of arLLM
[…]
@@ 6 Masked Fine-tuning of arLLM
Figure 3: An example of masked fine-tuning prompt. Random selection of text tokens are replaced by a [MASK] token. Highlighted tokens are used to compute the autoregressive loss.
@@
Inspired by the supremacy of dLLM in knowledge injection by fine-tuning, we attempt to adapt its advantages to arLLM. If an instruct arLLM is capable enough, one may prompt an arLLM to act like a dLLM. Specifically, given a masked document and an instruction to recover the original document, if the model has knowledge of the original document and the masked document contains sufficient cues to retrieve this knowledge, an instruct arLLM is supposed to respond with the correct original document. If the arLLM does not already have the knowledge of the original document, setting the ground truth document as the supervised fine-tuning target may implicitly teach the model that knowledge. We refer to this fine-tuning paradigm as “masked fine-tuning” of arLLM, and the resulting model as “masked arLLM”. Masked fine-tuning of arLLM, from a broader perspective, establishes a training objective similar to that of dLLMs, wherein the model learns to reconstruct the unmasked sequence from a masked input. Following the dLLM noise sampling strategy, we randomly replace sample tokens with a reserved special token during training, where the mask ratio $t$ is sampled from a uniform distribution $U(0.05,0.95)$. Note that each sample can be masked with a different mask ratio in different epochs. We evaluate the mask fine-tuned arLLM in the regular autoregressive way using the default chat template. The exact prompt used in the fine-tuning is provided in Figure 3 (more details in Appendix A.4), and the training objective formulation is:
@@
Overall, masked-finetuning of arLLM successfully inherits all the merits of the dLLM fine-tuning (Table 1, Figure 2). Masked arLLM surpasses arLLM fine-tuning in the pre-training style with a huge margin (Table 1) and achieves comparably high accuracy in both forward and backward question categories. Moreover, like dLLM, masked arLLM relies much less on paraphrases in the fine-tuning dataset to saturate the accuracy in most cases. The convergence rate of masked fine-tuning is also comparable to or in some cases faster than dLLMs (Figure 2, Table 4, Appendix Table 5), suggesting masked fine-tuning is data-efficient to achieve better down-stream QA tasks than traditional fine-tuning without costing additional training compute.
To confirm that the effectiveness of our masked fine-tuning stems from the objective itself and not merely a simple data augmentation effect (i.e., introducing varied input text prepended to the training sample), we conducted a control experiment, in which the masked document within the prompt was replaced entirely with random tokens (Appendix Figure 12). This substitution caused the accuracy of the masked fine-tuning to drop to the level of naive arLLM fine-tuning.
7 Effects of fine-tuning mask ratio
@@
Previous studies (zhu24a; zhu24b) claim that bidirectional BERT-like models struggle with even forward style knowledge extraction due to the mask loss, which causes the model to learn incorrect associations between tokens. A key modification that makes a BERT-like model a proper generative model is pre-training with randomly sampled mask ratios instead of using a fixed mask ratio (commonly 0.15 in BERT)(llada; bert). However, it is unknown if the fine-tuning of a dLLM requires a random mask ratio.
To investigate whether this necessity persists during post-training, we change the fine-tuning process of the dLLMs and masked arLLMs to use fixed mask ratios ($t$) instead of randomly sampling them during the training (Figure 4). Fine-tuning with some fixed mask ratios (0.75 and 0.5) can be as effective as the random mask ratio in knowledge injection. However, there is considerable performance variation across the choices of $t$. Conceptually, varying $t$ modulates task difficulty; a moderate $t$ places training in a “hard-but-not-impossible” regime that maximizes the per-example gradient signal. The effectiveness of some fixed mask ratios indicate that a dLLM only needs to sample over t to learn the demasking process during pre-training. Once this ability is acquired and not forgotten during fine-tuning, a fixed $t$ is sufficient for learning new knowledge.
Using a mask ratio of 0 in the masked fine-tuning of arLLM completely fails (black lines in Figure 4). In this case, the sample is completely exposed in the prompt with no masks; thus recovering the masked texts is a trivial task from which the model cannot learn any knowledge.
8 Masked supervised fine-tuning
To study whether the advantages of masked fine-tuning over regular fine-tuning of arLLM extend beyond knowledge injection QA tasks, we test this new paradigm with SFT to improve the model’s math capability. Different from previous sections in which fine-tuning samples are raw text containing new knowledge, the SFT fine-tuning samples contain QA pairs.
Models GSM8K MATH
@@ 8 Masked supervised fine-tuning
Table 3: Standard versus masked SFT across two math datasets (evaluated with 0-shot, pass@1). See full model names in section 3.
Similarly to the prompt construction used in Section 6 and Figure 3, to leverage the demasking training objective in arLLMs, we first transform QA pairs into a demasking task. Specifically, the user prompt consists of the question, a randomly masked answer, and an instruction to recover the full answer; the assistant response is the original full answer (See Appendix 16 for an example of the prompt). The loss formulation is the same as Eq. 2, but the value of $m$ is different in masked SFT, which is $1$ when a token is in the assistant response and its corresponding token in the user prompt is masked, and 0 elsewhere. We test masked SFT on two models, i.e. Llama-3.2-3B-Instruct and Qwen3-4B-Instruct-250, and two popular math datasets, i.e. GSM8K and MATH. Each dataset contains their own training set and testing set. Under all the conditions, masked SFT surpass the traditional SFT (Table 3). Experimental details including learning rate and epoch sweep are in the Appendix 16.
9 Computational cost comparison across methods
We use the Wiki dataset to comprehensively characterize the computational cost of different training methods tested in Table 1 covering data preparation, training, inference and convergence to peak accuracy. Specifically, we test 8b parameter arLLM (Llama-3.1-8b-instruct) and dLLM (Llada) models in the following 5 conditions: (1) arLLM + pretraining style finetuning + w/o paraphrases, (2) arLLM + pretraining style finetuning + w paraphrases, (3) arLLM + pretraining style finetuning + reverse training (golovneva2024reversetrainingnursereversal), (4) dLLM + pretraining style finetuning + w/o paraphrases, (5) arLLM + masked SFT style finetuning + w/o paraphrases (ours) (Table 4).
@@ 9 Computational cost comparison across methods
Data Preparation
Only condition (2) requires computational expensive data preparation to generate semantically identical and natural paraphrases of the original dataset. For example, in the Wiki dataset, paraphrasing 10 per sample costs 0.1M generation tokens of GPT-o3-mini or equivalent models. This cost scales further with larger original datasets and the requirements for the diversity of paraphrases. In contrast, all other methods require no or very simple data transformations during the training (sampling mask or reverse sequence) without additional compute.
Training
We compare training FLOPs, wall time and peak allocated memory per step. Theoretical FLOPs calculation follows formulation introduced in kaplan2020scalinglawsneurallanguage. Specifically, the training cost of a dense transformer is approximated as $2N$ FLOPs per token for the forward pass and $4N$ for backpropagation, in total $6N$ FLOPs per token, where $N$ denotes the model parameter size. For a step that processes $T$ tokens in total, the theoretical per-step cost is therefore $\mathrm{FLOPs/step}\;\approx\;6\,N\,T$ In our comparisons, we keep the model size fixed at N=8N{=}8B and use the same effective global batch size across all conditions, so the only varying factor is the average sequence length of training samples. We thus report the theoretical quantity as a factor of $T$. Since paraphrasing or reverse training only modify the information order within the training samples, condition (1)-(4) have comparable average sequence length $T$. Condition (5) requires presenting the model with the original sequence and its masked counterpart in SFT format (Figure 3) doubling the average sequence length to $2T+c$ where $c$ is the constant accounting for SFT instructions. We also empirically measure the average FLOPs, wall time and peak memory per step using PyTorch Profiler (pytorch_profiler_docs). Empirical FLOPs measurements agree with the theoretical calculations. Among all conditions, our proposed method condition (5) costs around twice training FLOPs and slightly larger wall time as well as peak allocated memory. Additionally, we report that the sampling the mask and constructing the masked fine-tuning prompt during training for condition (4) and (5) take negligible time (<3.5ms) and memory comparing to the forward and backward pass, thus causing no significant overhead.
Inference
For inference statistics, we only compare the theoretical inference FLOPs as empirical values highly depend on the generation length of a particular answer. Following the same calculation detailed in the above section, an inference step produces a forward pass through the dense transformer, leading to $\mathrm{FLOPs/step}\;\approx\;2\,N\,S$ where $S$ denotes the average generation sequence length. Only condition (4) dLLM needs inference FLOPs quadratic in $S$ as generation in dLLM cannot reuse the KV cache since it changes after each denosing step (ni2025difflm). All other conditions are linear in $S$.
Convergence
[…]
@@ 10 Discussion
We believe that such knowledge injection by fine-tuning will serve as a cornerstone for a self-evolving AI in the era of experience (silver2025welcome). Engineering a dynamic memory system for LLMs is a trending research field, as agentic LLMs need to learn and evolve from their experiences (zhang2025survey; chhikara2025mem0). Most current memory systems are based on external databases that store experiences and new knowledge as text. Such explicit textual memory has been successful due to the well-known in-context learning ability of LLMs. However, such memory systems have disadvantages: 1) a limited context window and degradation of performance with long context liu2023lost, 2) expensive computation due to long context, 3) difficulty in expressing implicit knowledge as text, such as knowledge of winning a chess game, and 4) the intrinsic limitation of using vector-based embeddings for retrieval (weller2025theoretical). Parametric memory (i.e., memorizing by changing the network weights) does not have the above issues; however, due to the complications of fine-tuning an LLM, parametric memory is much less popular in production settings (zhang2025survey). Furthermore, fine-tuning LLMs is shown to be inefficient in learning new factual knowledge (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization), and has to rely on gating to overcome catastrophic forgetting (pan2025memorization; wang2024wiserethinkingknowledgememory; yu2023meloenhancingmodelediting; hartvigsen2023aginggracelifelongmodel). Our study shows the feasibility of knowledge injection by fine-tuning through a mask recovery objective without additional paraphrases. These findings extend the known data efficiency of pre-training masked dLLMs (prabhudesai2025diffusion; ni2025difflm). The mask recovery objective uses a more flexible factorization, which can be seen as an implicit data augmentation. Therefore, it enables strong performance in both forward and backward style recalls. Furthermore, we show that such benefit is not exclusive to dLLM. The same objective can be reformulated into a supervised fine-tuning task for arLLM. We show that training arLLMs with this novel paradigm improves both knowledge injection and math SFT. This implies that one does not need to switch to a dLLM, but can use any of the existing arLLMs and still benefits from the data efficiency advantage.

### v3

@@ 4.1 Failures of arLLM knowledge injection
We first show that knowledge injection by fine-tuning in arLLMs heavily relies on paraphrases. This is known in previous studies (Berglund et al., 2023; Allen-Zhu and Li, 2025; Lin et al., 2024; Guo et al., 2024; Golovneva et al., 2024a). We demonstrate this observation on three datasets to set baselines for comparison with dLLM and our novel paradigm in the following sections.
@@
We fine-tune arLLMs on samples from each dataset using the pre-training format (i.e., without a chat template; using a chat template yields similar results, Figure 4 t=1.). Without paraphrases, backward accuracy on the NameDescription and Biography datasets is close to 0, while the forward accuracy of NameDescription N2D and Biography does not completely fail but is still poor (Table 1). Adding same-order paraphrases drastically raises forward accuracy close to 1, while backward accuracy remains close to 0. Paraphrases do not help backward accuracy in NameDescription and Biography datasets because the construction of these datasets does not change the semantic order of the sentences. The trend is similar in the Wiki dataset (Table 1 and 2). While the same-order paraphrases significantly increase forward accuracy, they only mildly increase backward accuracy. Using permute-order paraphrases increases both forward and backward accuracy, and the gap between them is smaller. Note that, due to the naturalness of this dataset, pre-fine-tuning accuracies are not as close to zero as in the other datasets (Table 1 and 2); nonetheless, they are sufficiently low to demonstrate the effectiveness of fine-tuning.
@@
These results suggest that, in arLLM fine-tuning, paraphrases significantly improve QA accuracy, but help backward questions only when the paraphrases change the information order in the original text to be more aligned with the backward style. Note that the accuracy difference between fine-tuning with paraphrases and without paraphrases is not due to different training steps; in both cases, we train the models with a sufficiently large number of epochs; the reported accuracy is taken from the best checkpoint during the training (Figure 2 and Appendix Figure 6).
@@
Figure 2: Training dynamics of arLLM (Llama 8B), dLLM (Llada), and masked arLLM (Llama 8B). For the NameDescription dataset, forward and backward accuracy are the average of N2D and D2N types. Paraphrases used in the Wiki dataset are the same-order paraphrases set. Due to the randomness of sampling the masks, we average across 4 random seeds for the dLLM and masked arLLM on NameDescription and Biography Datasets. Curves for each seed are shown in Appendix Figure 10-11.
4.2 Effectiveness of dLLM knowledge injection
Inspired by the known advantages of dLLMs described in the Introduction, we investigate whether they require paraphrases during fine-tuning to successfully handle both forward and backward QAs. We follow the original pretraining protocol (Nie et al., 2025b) to fine-tune LLaDA-8B-Instruct on the dataset samples using the loss defined in Eq. 1. On three datasets, the accuracy difference between fine-tuning with and without paraphrases is much smaller in the dLLM than in the arLLM (Table 1): dLLM without paraphrases can already achieve decent and comparable accuracies on both forward and backward questions; fine-tuning with paraphrases can further increase the accuracy by a small amount. Taken together, these results suggest that the dLLM is substantially less reliant on paraphrase augmentation during post-training and exhibits markedly reduced reversal-curse behavior in this setting. By plotting the test accuracy across the training steps (Figure 2), we observe that arLLM fine-tuned without paraphrases improves QA accuracy only in the beginning of training, then quickly decreases, indicating overfitting. The dLLM without paraphrases, on the other hand, does not show signs of overfitting. This finding echoes what has been found in comparing arLLMs and dLLMs in the pre-training phase (Prabhudesai et al., 2025; Ni et al., 2025).
One may expect that fine-tuning dLLM converges slower than arLLM, because learning any-order factorization requires seeing more than one way of the factorizations (i.e., samples masked in different ways) (Xue et al., 2025; Kim et al., 2025). However, we found that dLLM converges at least as fast as arLLM (Figure 2, Table 4, Appendix Table 5); in the Biography dataset, dLLM even converges faster than arLLM. This indicates that dLLM does not trade off better knowledge injection performance for more training compute; it requires the same or less training compute and fewer training samples but achieves better downstream performance.
4.3 Masked Fine-tuning of arLLM
@@
Figure 3: An example of masked fine-tuning prompt. A random selection of tokens is replaced by [MASK] token. Highlighted tokens are used to compute the autoregressive loss.
@@
Inspired by the supremacy of dLLM in knowledge injection by fine-tuning, we attempt to adapt its advantages to arLLM. If an instruct arLLM is capable enough, one may prompt an arLLM to act like a dLLM. Specifically, given a masked document and an instruction to recover the original document, if the model has knowledge of the original document and the masked document contains sufficient cues to retrieve this knowledge, an instruct arLLM should respond with the correct original document. If the arLLM does not already have the knowledge of the original document, setting the ground truth document as the supervised fine-tuning target may implicitly teach the model that knowledge. We refer to this fine-tuning paradigm as “masked fine-tuning” of arLLM, and the resulting model as “masked arLLM”. Masked fine-tuning of arLLM, from a broader perspective, establishes a training objective similar to that of dLLMs, wherein the model learns to reconstruct the unmasked sequence from a masked input. Following the dLLM noise sampling strategy, we randomly replace sample tokens with a reserved special token during training, where the mask ratio $t$ is sampled from a uniform distribution $U(0.05,0.95)$. Note that each sample can be masked with a different mask ratio in different epochs. We evaluate the masked fine-tuned arLLM in the regular autoregressive way using the default chat template. The exact prompt used in the fine-tuning is provided in Figure 3 (more details in Appendix A.4). The training objective is:
@@
Overall, masked fine-tuning of arLLM successfully inherits all the merits of the dLLM fine-tuning (Table 1, Figure 2). Masked arLLM surpasses arLLM fine-tuning in the pre-training style with a huge margin (Table 1) and achieves comparably high accuracy in both forward and backward question categories. Moreover, like dLLM, masked arLLM relies much less on paraphrases in the fine-tuning dataset to saturate the accuracy in most cases. The convergence rate of masked fine-tuning is also comparable to or in some cases faster than dLLMs (Figure 2, Table 4, Appendix Table 5), suggesting that masked fine-tuning is a more sample-efficient alternative for improving downstream QA than traditional fine-tuning, without requiring additional training compute.
To confirm that the effectiveness of our masked fine-tuning stems from the objective itself and not merely from a simple data augmentation effect (i.e., introducing varied input text prepended to the training sample target), we conducted a control experiment in which the masked document within the prompt was replaced entirely with random tokens (Appendix Figure 12). This substitution caused the accuracy of the masked fine-tuning to drop to the level of naive arLLM fine-tuning.
4.4 Effects of fine-tuning mask ratio
@@
Previous studies (Allen-Zhu and Li, 2024, 2025) claim that bidirectional BERT-like models struggle with even forward style knowledge extraction due to the mask loss, which causes the model to learn incorrect associations between tokens. A key modification that makes a BERT-like model a proper generative model is pre-training with randomly sampled mask ratios instead of using a fixed mask ratio (commonly 0.15 in BERT) (Nie et al., 2025b; Devlin et al., 2018). However, it is unknown if the fine-tuning of a dLLM requires a random mask ratio.
To investigate whether this necessity persists during post-training, we change the fine-tuning process of the dLLMs and masked arLLMs to use fixed mask ratios ($t$) instead of randomly sampling them during the training (Figure 4). Fine-tuning with some fixed mask ratios (0.75 and 0.5) can be as effective as the random mask ratio in knowledge injection. However, there is considerable performance variation across the choices of $t$. Conceptually, varying $t$ modulates task difficulty; a moderate $t$ places training in a “hard but not impossible” regime that maximizes the per-example gradient signal. The effectiveness of some fixed mask ratios indicates that a dLLM only needs to vary t to learn the demasking process during pre-training. Once this ability is acquired and not forgotten during fine-tuning, a fixed $t$ is sufficient for learning new knowledge.
Using a mask ratio of 0 in the masked fine-tuning of arLLM completely fails (black lines in Figure 4) because the reconstruction task becomes trivial (no missing tokens), yielding no learning signal. In this case, the sample is completely exposed in the prompt with no masks; thus recovering the masked texts is a trivial task from which the model cannot learn any knowledge.
4.5 Masked supervised fine-tuning
Table 3: Standard versus masked SFT across two math datasets (0-shot, pass@1). See full model names in Section 3.
Model GSM8K MATH
@@ 4.5 Masked supervised fine-tuning
To study whether the advantages of masked fine-tuning over regular fine-tuning of arLLM extend beyond factual knowledge injection QA tasks, we test this new paradigm with supervised fine-tuning (SFT) to improve the model’s procedural knowledge, i.e. math capability. Unlike previous sections in which fine-tuning samples are raw text containing new knowledge, the SFT fine-tuning samples contain QA pairs.
@@
Similarly to the prompt construction used in Section 4.3 and Figure 3, to leverage the demasking training objective in arLLMs, we first transform QA pairs into a demasking task. Specifically, the user prompt consists of the question, a randomly masked answer, and an instruction to recover the full answer; the assistant response is the original full answer (See Appendix 16 for an example of the prompt). The loss formulation is the same as Eq. 2, but the value of $m$ is different in masked SFT, which is $1$ when a token is in the assistant response and its corresponding token in the user prompt is masked, and 0 elsewhere. We test masked SFT on two models, i.e. Llama-3.2-3B-Instruct and Qwen3-4B-Instruct-2507, and two popular math datasets, i.e. GSM8K and MATH. Each dataset contains their own training set and testing set. Under all the conditions, masked SFT surpasses traditional SFT (Table 3). Experimental details including learning rate and epoch sweep are in the Appendix 16.
4.6 Computation cost comparison
We evaluate the end-to-end training-time cost of our masked autoregressive (masked-AR) fine-tuning against paraphrase-augmented autoregressive training and masked diffusion training, factoring in both (i) any extra compute required to prepare training data and (ii) the per-step training footprint (Table 4, analysis details in Appendix A.6). Paraphrase augmentation introduces a substantial pre-processing compute, since it requires generating multiple natural paraphrases per sample (e.g., 10× paraphrases cost 0.1M GPT-o3-mini generation tokens for Wiki, and it scales linearly with dataset size and desired diversity). In contrast, our masked-AR training uses only lightweight online transformations (mask sampling + prompt construction), which add negligible overhead compared to the forward/backward pass.
Our masked-AR method does increase per-step training compute because we present both the original sequence and its masked counterpart in an instruction-format example, effectively doubling the average sequence length (theoretical cost scales from T to 2T + c). Empirically, this shows up as  2× FLOPs/step (199.3 TFLOPs/step vs  92 TFLOPs/step for standard AR variants) and higher peak memory (51.1 GB vs 37.9 GB), with only a modest wall-time increase (2.76 s/step vs 2.32–2.84 s/step depending on baseline).
Crucially, this per-step overhead is offset by performance-relevant training efficiency: our method reaches the highest accuracy and exhibits a >2× faster convergence rate than the other methods under their best configurations, so the total training compute needed to reach peak accuracy is comparable—while avoiding any paraphrase-generation cost.
@@ 5 Discussion
We believe that knowledge injection via fine-tuning will serve as a cornerstone for self-evolving AI in the era of experience (Silver and Sutton, 2025). Engineering dynamic memory systems for LLMs is a trending research direction, as agentic LLMs need to learn and evolve from their experiences (Zhang et al., 2025; Chhikara et al., 2025). Most current memory systems are based on external databases that store experiences and new knowledge as text. Such explicit textual memory has been successful due to the well-known in-context learning ability of LLMs. However, these memory systems have several disadvantages: (1) a limited context window and degradation of performance with long contexts (Liu et al., 2023), (2) expensive computation due to long contexts, (3) difficulty in expressing implicit knowledge as text, such as knowledge of how to win a chess game, and (4) intrinsic limitations of vector-based embeddings for retrieval (Weller et al., 2025). Parametric memory (i.e., memorizing by changing the network weights) does not have these issues; however, due to the complications of fine-tuning an LLM, parametric memory is much less popular in production settings (Zhang et al., 2025).
Inspired by recent dLLM studies, we demonstrate that knowledge injection via fine-tuning can be substantially improved through a demasking objective for arLLMs. Overall, our results suggest that the demasking objective provides a simple, sample-efficient route to reliable parametric knowledge updates in autoregressive LLMs. We hope future work scales this paradigm to continual, real-world memory and agent learning settings.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning by improving the sample efficiency and generalization of knowledge injection during fine-tuning of autoregressive language models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.